40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 10 15:05:32 2014, Jenne, Update, Electronics, Xend QPD whitening board modified already IMG_1776.JPG
    Reply  Thu Dec 11 16:42:12 2014, Jenne, Update, Electronics, Xend QPD whitening board plan QPDwhiteningModification_11Dec2014.pdf
       Reply  Fri Dec 12 19:54:21 2014, Jenne, Update, Electronics, Xend QPD whitening board modified QPDwhiteningModification_XtransCompleted_12Dec2014.pdfIMG_1779.JPGQPD_Xtrans_Ytrans_12Dec2014.pdfQPD_Xtrans_Ytrans_DarkVsLock_12Dec2014.pdf
          Reply  Sat Dec 13 00:35:11 2014, rana, Update, Electronics, Xend QPD whitening board modified 
          Reply  Mon Dec 15 22:30:50 2014, Jenne, Update, Electronics, Yend QPD modified 
             Reply  Mon Dec 15 22:45:59 2014, Jenne, Update, Electronics, Yend QPD modified QPD_Ytrans_oscillating_15Dec2014.pdfQPD_IOPchannels_Ytrans_oscillating_15Dec2014.pdf
                Reply  Fri Feb 13 14:27:30 2015, ericq, Update, Electronics, Second QPD Whitening Switch enabled 
                   Reply  Fri Feb 13 14:59:13 2015, ericq, Update, Electronics, Second QPD Whitening Switch enabled 
                      Reply  Fri Feb 13 18:56:44 2015, rana, Update, Electronics, Second QPD Whitening Switch enabled 
                         Reply  Thu Feb 19 04:09:21 2015, ericq, Update, Electronics, Second QPD Whitening Switch enabled 
                   Reply  Fri Feb 13 17:07:51 2015, Jenne, Update, Electronics, Second QPD Whitening Switch enabled 
Message ID: 11022     Entry time: Fri Feb 13 14:27:30 2015     In reply to: 10801     Reply to this: 11023   11024
Author: ericq 
Type: Update 
Category: Electronics 
Subject: Second QPD Whitening Switch enabled 

I have re-enabled the second whitening stage switching on each quadrant of each end's QPD whitening board, to try and avoid saturations at full power. Looking at the spectra while single arm locked, I confirmed that the FM2 whitening switch works as expected. FM1 should be left on, as it is still hard-wired to whiten. 

The oscillations in the Y QPD still exist. Jenne is updating the schematics on the DCC.

ELOG V3.1.3-